OpenStudy (anonymous):

lim ((1/c+h)-(1/c))/h h → 0

OpenStudy (anonymous):

\[\lim_{h \rightarrow 1} \frac{ ((\frac{ 1 }{ c+h })-(\frac{ 1 }{ c })) }{ h }\] To make the equation more clear.

OpenStudy (anonymous):

sorry, as h-->0

OpenStudy (p0sitr0n):

put over common denominator c-c+h ______ = h/(c+h)h=1/(c+h) / h= 1/(c+h)h= 1/(c+0)0 = 1/0=+infinity (c+h)h

OpenStudy (anonymous):

\[\lim_{h \to 1} \frac{ \dfrac{ 1 }{ c+h }-\dfrac{ 1 }{ c } }{ h }\\ \lim_{h \to 1} \frac{ \dfrac{ c }{ c(c+h) }-\dfrac{ c+h }{ c(c+h) } }{ h }\\ \lim_{h \to 1} \frac{ \dfrac{ c -c-h}{ c(c+h) }}{ h }\\ -\lim_{h \to 1} \frac{ \dfrac{ h}{ c(c+h) }}{ h }\\ -\lim_{h \to 1} \dfrac{1}{ c(c+h) }=\cdots \]